OpenStudy (anonymous):

Identify the conjugate of: 4+√7 1)√7-4 2)7+√4 3)4+√7 4)4-√7

OpenStudy (anonymous):

4) 4 - √7

myininaya (myininaya):

4 or 1 is the answer addition is commutative \[4+\sqrt{7}\]=\[\sqrt{7}+4\] the conjugate of \[4+\sqrt{7} is 4-\sqrt{7}\] the conjugate of \[\sqrt{7}+4is \sqrt{7}-4=-4+\sqrt{7}=\sqrt{7}-4\]

myininaya (myininaya):

since \[4+\sqrt{7}=\sqrt{7}+4\] then you can say the conjugate is either \[4-\sqrt{7}\] or \[\sqrt{7}-4\]